OpenStudy (anonymous):

tan (x) csc (x) I need to match the above trig expression to one of the following: sec (x), -1, cot (x), 1, -tan (x), or sin (x) An explanation would also be great! I'm pretty sure I need to use an Identity, but I'm not sure which one.

OpenStudy (anonymous):

do you know what is csc(x) ?

OpenStudy (anonymous):

csc (x) is cosecant, the reciprocal of sin (x)

OpenStudy (anonymous):

good and tan(x) ?

OpenStudy (anonymous):

sin(x)/cos(x), right?

OpenStudy (anonymous):

right so what would be the multiplication of them ?

OpenStudy (anonymous):

I'm not sure I follow the question, do you mean the product of sin and cos, or the product of tan and csc?

OpenStudy (anonymous):

we know that tan(x) = sinx(x)/cos(x) and csc(x) = 1/sin(x) we need to find what is tan(x) * csc(x) = (sin(x)/cos(x)) * (1/sin(x))

OpenStudy (anonymous):

do you know the answer now ?

OpenStudy (anonymous):

so essentially all I need to do is simplify (sin(x)/cos(x)) * (1/sin(x))? If so, I should be able to work that out

OpenStudy (anonymous):

yes .. in fact there is nothing much to do in this question you only had to write tan(x) and csc(x) as the basic functions sin(x),cos(x) :)

OpenStudy (anonymous):

Wow. It's seems so simple now! The answer is Sin(x), yes?

OpenStudy (anonymous):

\[\frac{ \sin(x) }{ \cos(x) } * \frac{ 1 }{ \sin(x) }\]

OpenStudy (anonymous):

are you sure ?

OpenStudy (anonymous):

I was taught never to cancel out, so I need to cross multiply. (Sin^2(x) cos(x))/(cos(x) sin(x)) The cosines divide out to 1, leaving me with Sin^2(x)/Sin(x), which then divides out to Sin(x)

OpenStudy (anonymous):

Have I done anything wrong?

OpenStudy (anonymous):

yes seems like you treat it as an equation : \[\frac{ sinx }{ cosx } = \frac{ 1 }{ sinx }\] ? or i dont even understand what you did there

OpenStudy (anonymous):

Cross multiply both Sin(x) for Sin^2(x), then cross multiply Cos(x) and 1, which is just Cos(x) After that, you multiply the denominators, which stay on the bottom, so that gives you Cos(x) Sin(x) The numerator comes out to be Sin^2(x) Cos(x) and the denominator comes out to be Cos(x) Sin(x), ergo (Sin^2(x) Cos(x))/(Cos(x) Sin(x))

OpenStudy (anonymous):

but why do you think of cross multiplication you do this when you have an equation .. this is an expression!!

OpenStudy (anonymous):

We're multiplying two fractions though, so cross multiply, right?

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

tell me please my friend : \[\frac{ 1 }{ 2 } * \frac{ 6 }{ 1 }\]

OpenStudy (anonymous):

what is the result ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right .. no cross multiplication here agree?

OpenStudy (anonymous):

OH, I see what you mean now!

OpenStudy (anonymous):

I just did a quick redo without cross multiplying and got Sec(x)

OpenStudy (anonymous):

there is a point anyway .. whenever you cancel out something you have to make sure it isnt zero .. so in fact the solution holds only if sin(x) isnt zero which is obvious cause csc(x) wont exist if so

OpenStudy (anonymous):

yes.. the answer is sec(x) :)

OpenStudy (anonymous):

Thank you for walking me through this! Much appreciated!

OpenStudy (anonymous):

yw :)